OpenStudy (anonymous):

Sove by Elinination 3x+3y=1 and 4x+6y=2

OpenStudy (anonymous):

multiply the 1st equation by 2 and subtract 2nd equation from 1st equation,y term will be canceled and you will get the x value.then plug x value into one of the equation to get y value.

OpenStudy (anonymous):

I have the answer 6

OpenStudy (anonymous):

I am not getting any of this

OpenStudy (anonymous):

next question is find the indicated outputs for F(x)-3xsquare-2x

OpenStudy (anonymous):

multiply the 1st equation by 2 2(3x+3y)=2(1)-----> 6x+6y=2 now,subtract 2nd equation from 1st equation 6x+6y=2 -4x-6y=-2 ----------- 2x=0 x=0 now plug x into one of the equation 3(0)+3y=1 3y=1 y=1/3 so,x=0 and y=1/3

OpenStudy (anonymous):

Thanks you make it easier I have no knowledge on this and it is required ugh

OpenStudy (anonymous):

we multiply equation i with 2 and u will get equation 6x+6y=2........................i then subtract equation 4x +6y=2 from equation i u get 2x=0 so x=0 putting value of x in any equation u will get y=1/3

OpenStudy (anonymous):

Sove by Elinination 3x+3y=1 and 4x+6y=2 2*(3x +3y = 1) is 6x +6y =2 Subtract equation(2) 4x + 6y =2 we get 2x =0 , x =0 Plug in x =0 in the first equation 3x +3y = 1, 0 +3y = 1 3y = 1 Divide by 3 y = 1/3